OpenStudy (anonymous):

Solve for x: x-1/9 (over 9!) =2 Options... 18 17 12 19 Thank you:) I'm so stuck

OpenStudy (uri):

x-1/9 =2 x=2 x 9 -1 x=18-1 x=17.

OpenStudy (anonymous):

Thank you!